A partical of mass m is hung by two light strings.the ends A and B are held by hands. The strings OA and OB make angle q with the vertical. Find the values of T and T' in terms of m and q
1
Expert's answer
2021-08-22T15:47:00-0400

Gives

Angle=q




Daigram "\theta=q"

Lami theorem

"\frac{T}{sin(90+q)}=\frac{T}{sin(90+q)}=\frac{mg}{sin(180-2q)}" Second and third part we can written as

"T=\frac{mg\times sin(90+q)}{sin(180-2q)}"

"T=\frac{mg\times cos(q)}{sin(2q)}"

"T=\frac{mg\times cos(q)}{2sin(q)cos(q)}"

"T=\frac{mg}{2sin (q)}"